OpenStudy (anonymous):

write f(x) = x^3 +4x^2+4x+16 as a product of linear factors i think the answer is x =(x-4)(x+2)^

OpenStudy (anonymous):

that not one of the answer choice in the question but i appreciate the help

OpenStudy (alekos):

what are the choices?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

A x =(x-4)(x+2)^2 B x= (X-4)(X-2)^2 c x =(x+4)(x-4)(x+) d x=(x+4)(x+2i)(x-i) e x= (x-4)^2(x-2i)

OpenStudy (alekos):

should be (x+4)(x+2i)(x-2i)

OpenStudy (loser66):

I prove A) (x-4) (x+2)^2 = (x-4)(x^2+4x+4) = x^3 +4x^2+4x-4x^2-16x-16= x^3-12x-16 not the original problem, reject B)(x-4)(x-2)^2= (x-4)(x^2-4x+4) = x^3 -4x^2+4x-4x^2+16x-16= x^3-8x^2+20x-16 not the original problem, reject C) (x+4)(x-4)(x+? ) no matter what the last one is, it 's still wrong. because the first 2 terms give you x^2 -16 and the last term of the original one is +16, not -16, done D) (x+4) (x+2i)(x-i) never be a correct answer, because for the imaginary root, it is always a pair of it, it is never be 2i and i E) (x-4)^2(x-2i) the same logic with the above

OpenStudy (anonymous):

thank you both for the answer and explaining it

OpenStudy (loser66):

even though the option D is (x+4)(x+2i)(x-2i) = (x+4) (x^2-(2i)^2) = (x+4) (x^2 -4) = x^3-4x+4x^2 -16 reject still.

OpenStudy (loser66):

so, I shouldn't sue the calculator company, right? hahahaa

OpenStudy (loser66):

You don't have the correct options or YOU POST THE WRONG PROBLEM hehehe

OpenStudy (anonymous):

im going with C it has to be right

OpenStudy (loser66):

it's up to you

OpenStudy (alekos):

I think option d should have a -2i in the last bracketed term (X+4)(X+2i)(x-2i) = (x+4)(x^2+4)